# The decomposition of primes in nonabelian extensions of Heisenberg type and an analogue of Euler’s criterion

### v1

@@ 1. Introduction
for distinct odd primes $p$ and $q$ such that 44 divides p1p-1. In the language of class field theory, it relates the Artin symbols at $q$ in two abelian extensions $\mathbb{Q}(\sqrt{p})$ and $\mathbb{Q}(\zeta_{p})$ of $\mathbb{Q}$, where $\zeta_{m}$ denotes a primitive $m$-th root of unity for $m\geq 1$.
@@
For our purpose, we will take $k=\mathbb{F}_{p}$ for a prime pp. Since we need a primitive $\ell$-th root of unity, we always assume $\ell\mid p-1$. Our main result concerns the splitting of the prime ideal (ta)(t-a) for $a\in\mathbb{F}_{p}-\{0,1\}$ in the mod $\ell$ Heisenberg extension $\mathcal{R}^{(\ell)}/\mathbb{F}_{p}(t)$. To state it, we introduce certain polynomials that will play the role of $x^{(p-1)/2}$ for $\mathbb{Q}(\sqrt{p})$. Put $\epsilon_{\ell,n}(t)=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}$. Note that $\epsilon_{\ell,0}(t)$ is literally $\epsilon_{\ell}(t)$ which is a generator of $\mathcal{R}^{(\ell)}$. Define the polynomial
@@ 3. $\ell=2$
All elements in the mod 22 Heisenberg group have order of at most 44.
@@
An arbitrary $A\in\mathcal{H}(\mathbb{F}_{2})$ can be written as $A=I+B$ where II is the $3\times 3$ identity matrix and $B$ is a strict upper-triangular matrix. Then, $A^{4}=(I+B)^{4}=I+B^{4}=I$ gives the required result. ∎
@@
Since the Frobenius conjugacy class that corresponds to (ta)(t-a) has order of at most 44, the number of prime ideals above (ta)(t-a) is at least 22 by Proposition 1. ∎
@@
Recall from (2) that $A_{2}(a)=\frac{1}{2}\left((1-\sqrt{a})^{\frac{p-1}{2}}+(1+\sqrt{a})^{\frac{p-1}{2}}\right)$. So, if $\genfrac{(}{)}{}{}{1-x}{p}=1$, the value of the above polynomial is 11, and if $\genfrac{(}{)}{}{}{1-x}{p}=-1$, the value of the above polynomial is 1-1. Finally, since $x^{2}=a$ in $\mathbb{F}_{p}$ and $\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{1-p}{2}}\right)$ only has even degree terms, we obtain our desired result. ∎
@@
Hence, if $\genfrac{(}{)}{}{}{a}{p}=\genfrac{(}{)}{}{}{1-a}{p}=1$, $x_{\frac{p-1}{2}}=A_{2}(a)$ is 11 or 1-1 which is determined by $\genfrac{(}{)}{}{}{1-\sqrt{a}}{p}$ as we argued above. If $\genfrac{(}{)}{}{}{a}{p}=1$ and $\genfrac{(}{)}{}{}{1-a}{p}=-1$, then $x_{\frac{p-1}{2}}^{2}=0$, which is equivalent to $A_{2}(a)=0$.
@@ 4. $\ell\geq 3$
An arbitrary $A\in\mathcal{H}(\mathbb{F}_{\ell})$ can be written as $A=I+B$ where II is the $3\times 3$ identity matrix and $B$ is a strict upper-triangular matrix. Then, $A^{\ell}=(I+B)^{\ell}=I+B^{\ell}=I$ since $\ell\geq 3$. ∎
@@
From the formula for determinant of the Vandermonde matrix, $\det B=\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)$. Now, consider the series of row operations $\Set{R_{i}}$ that transforms $B$ into diagonal matrix $B^{\prime}$. Then, the product of diagonal elements is given by $\det B$. Now, apply the same row action on AA considering it as a $\ell\times\ell$ matrix with $n\times n$ block matrix as its entry. Then, we get the block diagonal matrix with diagonal entries $B^{\prime}_{ii}A_{i}$. Hence, $\det A=\prod_{i}\det(B_{ii}^{\prime}A_{i})=(\det B)^{n}\times\prod_{i}\det(A_{i})$

### v2

@@ 1. Introduction
for distinct odd primes $p$ and $q$ such that $4$ divides p1p-1. In the language of class field theory, it relates the Artin symbols at $q$ in two abelian extensions $\mathbb{Q}(\sqrt{p})$ and $\mathbb{Q}(\zeta_{p})$ of $\mathbb{Q}$, where $\zeta_{m}$ denotes a primitive $m$-th root of unity for $m\geq 1$.
@@
For our purpose, we will take $k=\mathbb{F}_{p}$ for a prime $p$. Since we need a primitive $\ell$-th root of unity, we always assume $\ell\mid p-1$. Our main result concerns the splitting of the prime ideal (ta)(t-a) for $a\in\mathbb{F}_{p}-\{0,1\}$ in the mod $\ell$ Heisenberg extension $\mathcal{R}^{(\ell)}/\mathbb{F}_{p}(t)$. To state it, we introduce certain polynomials that will play the role of $x^{(p-1)/2}$ for $\mathbb{Q}(\sqrt{p})$. Put $\epsilon_{\ell,n}(t)=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}$. Note that $\epsilon_{\ell,0}(t)$ is literally $\epsilon_{\ell}(t)$ which is a generator of $\mathcal{R}^{(\ell)}$. Define the polynomial
@@ 3. $\ell=2$
All elements in the mod $2$ Heisenberg group have order of at most $4$.
@@
An arbitrary $A\in\mathcal{H}(\mathbb{F}_{2})$ can be written as $A=I+B$ where $I$ is the $3\times 3$ identity matrix and $B$ is a strict upper-triangular matrix. Then, $A^{4}=(I+B)^{4}=I+B^{4}=I$ gives the required result. ∎
@@
Since the Frobenius conjugacy class that corresponds to (ta)(t-a) has order of at most $4$, the number of prime ideals above (ta)(t-a) is at least $2$ by Proposition 1. ∎
@@
Recall from (2) that $A_{2}(a)=\frac{1}{2}\left((1-\sqrt{a})^{\frac{p-1}{2}}+(1+\sqrt{a})^{\frac{p-1}{2}}\right)$. So, if $\genfrac{(}{)}{}{}{1-x}{p}=1$, the value of the above polynomial is $1$, and if $\genfrac{(}{)}{}{}{1-x}{p}=-1$, the value of the above polynomial is 1-1. Finally, since $x^{2}=a$ in $\mathbb{F}_{p}$ and $\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{1-p}{2}}\right)$ only has even degree terms, we obtain our desired result. ∎
@@
Hence, if $\genfrac{(}{)}{}{}{a}{p}=\genfrac{(}{)}{}{}{1-a}{p}=1$, $x_{\frac{p-1}{2}}=A_{2}(a)$ is $1$ or 1-1 which is determined by $\genfrac{(}{)}{}{}{1-\sqrt{a}}{p}$ as we argued above. If $\genfrac{(}{)}{}{}{a}{p}=1$ and $\genfrac{(}{)}{}{}{1-a}{p}=-1$, then $x_{\frac{p-1}{2}}^{2}=0$, which is equivalent to $A_{2}(a)=0$.
@@ 4. $\ell\geq 3$
An arbitrary $A\in\mathcal{H}(\mathbb{F}_{\ell})$ can be written as $A=I+B$ where $I$ is the $3\times 3$ identity matrix and $B$ is a strict upper-triangular matrix. Then, $A^{\ell}=(I+B)^{\ell}=I+B^{\ell}=I$ since $\ell\geq 3$. ∎
@@
From the formula for determinant of the Vandermonde matrix, $\det B=\left(\prod_{i\neq j}(\zeta_{\ell}^{i}-\zeta_{\ell}^{j})\right)$. Now, consider the series of row operations $\Set{R_{i}}$ that transforms $B$ into diagonal matrix $B^{\prime}$. Then, the product of diagonal elements is given by $\det B$. Now, apply the same row action on $A$ considering it as a $\ell\times\ell$ matrix with $n\times n$ block matrix as its entry. Then, we get the block diagonal matrix with diagonal entries $B^{\prime}_{ii}A_{i}$. Hence, $\det A=\prod_{i}\det(B_{ii}^{\prime}A_{i})=(\det B)^{n}\times\prod_{i}\det(A_{i})$